Years Present Value 900 2,591 34,105 33,800 Interest Rate 12 % 10 Future Value $ 1,755 4,350 393,620 217,868 22
Answer #1
FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
a) FV/ PV = (1 + r )^n
1755 / 900 = (1 + 12%)^n
1.95 = 1.12^n
n = 5.89 years
b) FV/ PV = (1 + r )^n
4350 / 2591 = (1 + 10%)^n
1.6788884600= 1.10^n
n = 5.44 years
c) FV/ PV = (1 + r )^n
393620 / 34105 = (1 + 15%)^n
11.541416214 = 1.15^n
n = 17.50 years
d) FV/ PV = (1 + r )^n
217868 / 33800 = (1 + 22%)^n
6.4457988165 = 1.22^n
n = 9.37 years